### v3

# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

Abstract
We propose a general approach to optimize noise schedules for training and sampling in diffusion models. Our approach optimizes the noise schedules to ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process. Any distance metric for measuring the probability-distributional change is applicable to our approach, and we introduce three distance metrics. We evaluated the effectiveness of our approach on unconditional and class-conditional image-generation tasks using the LSUN (Horse, Bedroom, Church), ImageNet, FFHQ, and CIFAR10 datasets. Through extensive experiments, we confirmed that our approach broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations ranging from 5 to 250. Notably, by using our approach for optimizing both training and sampling schedules, we achieved a state-of-the-art FID score of 2.03 without sacrificing mode coverage on LSUN Horse 256 $\times$ 256.
1 Introduction
Diffusion models are a type of probabilistic generative model that have gained significant attention for their expressive ability to generate high-quality samples across various domains, such as image generation image_generation1 ; image_generation2 ; image_generation3 , image super-resolution image_super_resolution1 ; image_super_resolution2 , video synthesis video_generation1 ; video_generation2 , and audio generation audio_generation1 ; audio_generation2 .
Diffusion models were originally proposed in the previous study nonequiliburium_thermodynamics , and denoising diffusion probabilistic models (DDPMs) ddpm_original demonstrate that diffusion models can achieve state-of-the-art performance in image generation. In diffusion models, two diffusion processes are defined using Markov chains, referred to as the forward process and reverse process. In the forward process, data are diffused by gradually adding Gaussian noise, until the probability distribution of diffused data becomes a Gaussian distribution. In the reverse process, data are generated from Gaussian noise by tracing the forward process in the reverse direction. Noise levels added to data at each step of the diffusion processes are hyperparameters and referred to as the noise schedule. We do not necessarily have to use a same noise schedule for training and sampling, and recent studies edm ; vdmpp have shown that using different noise schedules can lead to improved performance. Although diffusion models tend to perform well in terms of sample fidelity and mode coverage, they face the challenge of longer generation times due to the iterative function evaluations during sampling diffusion_gans .
To improve the performance of diffusion models, we propose general approach for optimizing training and sampling schedules that ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process. We call our approach constant rate scheduling (CRS). As previously stated, the reverse process needs to accurately trace the forward process in the reverse direction to restore the data from pure noise. This motivates us to determine noise schedules that can enhance the traceability of the forward process. Figure 1 illustrates a toy example of diffused data distributions with three data points in one-dimensional data space (see Appendix A for details). Here, $\alpha$ represents the decay rate of data at each step of the diffusion process. In this example, the diffusion process can be divided into three regions. The first is an area where the probability distributions barely change, and the noise level can be changed rapidly ($\alpha\lesssim 0.6$). The second is an area where three modes corresponding to the data points emerge, and careful adjustment of the noise level is necessary to enhance the mode coverage ($0.6\lesssim\alpha\lesssim 0.97$). The third is an area where the peaks of each mode become distinct, and the noise level should be changed slowly for sample fidelity ($\alpha\gtrsim 0.97$). As evident from this example, the diffusion process exhibits regions with rapid and slow changes in the probability distributions. We assume that the probability-distributional change of diffused data represents the traceability of the diffusion process and determines the noise schedule to ensure a constant rate of change in the probability distributions.
CRS can be considered general from the following perspectives: 1) CRS can optimize both the training and sampling schedules, 2) any distance metric for measuring probability-distributional change is applicable, allowing the introduction of domain specific distance metrics and the use of different noise schedules for training and sampling, 3) any noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
Our contributions are summarized as follows:
  1. 1.
    We propose a general approach to optimize noise schedules for both training and sampling. We can use any distance metric for measuring probability-distributional change and introduce three distance metrics. One is specialized for image generation, while the other two are expected to be applicable across a wide range of domains beyond image generation.
  2. 2.
    Through extensive experiments, we confirm that CRS broadly improves the performance of image generation in pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations (NFEs) from 5 to 250. CRS contributes to the performance improvement at both training and sampling, achieving a state-of-the-art Fréchet inception distance (FID) fid_score score of 2.03 on LSUN Horse 256 $\times$ 256 dataset. CRS is fundamentally different from conventional approaches, such as align your steps (AYS) noise_ays and geometry-inspired time scheduling (GITS) noise_gits , in that it can optimize the training schedule as well.
Figure 1: Toy example of diffused data distributions with three data points in one-dimensional data space. Probability distributions barely change when $\alpha\lesssim 0.6$, and we can rapidly change noise level. Three modes corresponding to data points emerge when $0.6\lesssim\alpha\lesssim 0.97$, and noise level should be changed slowly for mode coverage. Three modes become distinct when $\alpha\gtrsim 0.97$, requiring careful control of noise level for sample fidelity.
2 Related Work
Recent studies regarding fast sampling can be mainly categorized into the following three strategies.
Samplers: This strategy devises update rules of diffused data in the reverse process. Denoising diffusion implicit models (DDIMs) ddim_original introduce a non-Markov chain to generalize DDPMs and propose a deterministic sampler. In the continuous-time diffusion models, the forward and reverse processes are formulated using a stochastic differential equation (SDE) and ordinary differential equation (ODE) score_matching3 . Numerous studies have been conducted to develop efficient solvers tailored for the reverse SDE and ODE in diffusion models sampler_adaptive_step ; sampler_pndm ; sampler_dpm_solver ; sampler_dpmpp ; sampler_ipndm ; sampler_unipc ; sampler_dpm-v3 ; sampler_amed .
Noise schedules: Noise schedules significantly affect the performance of diffusion models noise_effect , and many noise schedules have been proposed ddpm_original ; improved_diffusion ; noise_effect ; shifted_cosine . However, the functional form in these studies, such as the linear ddim_original and cosine improved_diffusion functions, is determined somewhat arbitrarily. Recently, systematic optimization of noise schedules have begun to be actively studied noise_gits ; noise_jys . The most relevant approach to CRS is AYS noise_ays , which determines a noise schedule for sampling to minimize the discretization errors when solving the reverse SDE. However, noise schedules aimed at improving the training efficiency have been largely overlooked compared with the sampling efficiency, though several studies, such as elucidating diffusion models (EDMs) edm and variational diffusion models++ (VDM++) vdmpp , have investigated them.
Distillations: The distillation-based acceleration is one of the most important direction for fast sampling distill_org1 ; distill_org2 ; distill_consistency ; distill_dmd ; distill_dmd2 ; distill_s4s . CRS is expected to provide better teacher models for distillation methods because it can optimize the training schedule as well.
3 Background
We briefly introduce DDPMs.
In DDPMs ddpm_original , the forward process, which starts from a data point $\bm{x}_{0}\sim q(\bm{x})$, is defined as
$q(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})\right)q(\bm{x}_{0})\ \ \mathrm{and}\ \ q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\beta_{t}\bm{x}_{t-1},\delta_{t}^{2}\bm{I}),$ (1)
where $\beta_{t}$ is the pre-defined decay rate of the data, $\delta_{t}$ represents the noise strength added to $\bm{x}_{t-1}$, and $T$ is the number of timesteps in the forward process. It is straightforward to verify that
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I}),$ (2)
where $\alpha_{t}=\beta_{t}\alpha_{t-1}$ and $\sigma_{t}^{2}=\delta_{t}^{2}+\beta_{t}^{2}\sigma_{t-1}^{2}$. We focus on the variance-preserving process, where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ and $\beta_{t}^{2}+\delta_{t}^{2}=1$ are satisfied.
The reverse process starts from Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. In the limit of $T\to\infty$, the probability distribution of the reverse process is well approximated by the Gaussian distribution ddpm_approx :
$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\nu_{t}^{2}\bm{I}).$ (3)
Using this approximation, the noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},t)$:
$\bm{\mu}_{\theta}(\bm{x},t)=\frac{1}{\beta_{t}}\left(\bm{x}_{t}-\frac{\delta_{t}^{2}}{\sigma_{t}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)\right),$ (4)
is optimized using a simplified loss function derived from the variational lower bound.
4 General Approach for Optimizing Noise Schedule
We explain our motivation and present CRS in more detail.
4.1 Motivation
As stated in Sec. 3, when $T$ is sufficiently large, it is a good approximation to assume the probability distribution of the reverse process $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ as a Gaussian distribution, as expressed in Eq. (3). In other words, the distributional change between adjacent timesteps must be small enough to be well approximated by a Gaussian distribution. This observation motivates us to minimize the maximum distributional change in the diffusion process for holding the approximation in Eq. (3) throughout the diffusion process.
We can understand the importance of the probability-distributional change in the diffusion process from the perspective of noise conditional score networks (NCSNs) score_matching1 ; score_matching2 ; score_matching3 , which are also the origin of diffusion models. According to the manifold hypothesis, natural data typically lie on a low-dimensional manifold embedded within a high-dimensional data space. Based on this property, there exist two main challenges in generating samples using only the score of the data distribution $\nabla_{\bm{x}}\log q(\bm{x})$ starting from a random initialization $\bm{x}_{\mathrm{rnd}}$ score_matching1 . First, $\bm{x}_{\mathrm{rnd}}$ is likely to be located in regions with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, where it is difficult to learn the score accurately. Second, even if the score can be learned correctly, the score tends to be small in the region with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, leading to slow mixing. To address these challenges, NCSNs interpolate between the data distribution and a Gaussian distribution, the support of which is the whole space. If the distributional change between adjacent timesteps is sufficiently small during the interpolation process, a sample generated at a certain timestep is likely to fall into a region where the probability density of the adjacent timestep is high. Therefore, by minimizing the maximum distributional change between adjacent timesteps, it becomes possible to generate samples using the reliable score at all timesteps of the interpolation process.
4.2 Constant Rate Scheduling
As described above, our criterion to optimize noise schedules is to minimize the maximum distance between consecutive probability distributions in the diffusion process. Using the distance $D(t,t+\Delta t)$ between the probability distributions of diffused data $p(\bm{x}_{t})$ and $p(\bm{x}_{t+\Delta t})$, our criterion is expressed as
$\underset{\alpha(t)}{\operatorname{argmin}}\left(\max_{t}D(t,t+\Delta t)\right).$ (5)
This optimization problem has a straightforward solution as follows:
$D(t,t+\Delta t)=\frac{D(t,t+\Delta t)-D(t,t)}{\Delta t}\Delta t\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=\mathrm{const.},$ (6)
where we use $D(t,t)=0$.
To explicitly show the dependence on the noise schedule, we rewrite $D(t,t^{\prime})$ as $\tilde{D}(\alpha(t),\alpha(t^{\prime}))\equiv\tilde{D}(\alpha,\alpha^{\prime})$. Here, $\alpha(t)$ is the noise-schedule function to be optimized. This results in
$D(t,t+\Delta t)\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=-v(\alpha)\frac{d\alpha(t)}{dt}\Delta t=\mathrm{const.},$ (7)
where $v(\alpha)=\left.\frac{\partial\tilde{D}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}$ represents the rate of change in the probability distributions of diffused data. Therefore, the noise-schedule function is determined to satisfy $-\frac{d\alpha(t)}{dt}\propto v(\alpha)^{-1}$. The proportionality coefficient can be calculated using the boundary conditions: $\alpha(0)=\alpha_{\max}$ and $\alpha(1)=\alpha_{\min}$. Finally, we obtain the noise-schedule function as follows:
[…]
Here, we introduce a hyperparameter $\xi>0$ to control the dependence of $\alpha(t)$ on $v(\alpha)$. As $\xi$ is increased, more timesteps are allocated to the region where $v(\alpha)$ is large.
Similar to NCSN (score_matching1, ) inspired by simulated annealing (sa_original, ), our formulation draws inspiration from the adiabatic theorem (adiabatic_theorem, ) in quantum annealing (qa_original, ). Quantum annealing is an algorithm for solving combinatorial optimization problems, where the target cost function to be optimized is interpolated with an initial cost function with a trivial optimal solution. The initial state of the optimization is set to the optimal solution of the initial cost function. By tracing the optimal solutions of the instantaneous cost functions during the interpolation, the optimal solution of the target cost function is then obtained. In quantum annealing, accurately tracing the optimal solution of an instantaneous cost function is crucial, and the adiabatic theorem derives the traceability of the optimal solution at each step. Let $\alpha(t)$ be the weight of the target cost function and $1-\alpha(t)$ be the weight of the initial cost function. In this case, the inverse traceability can be expressed in the form of $\frac{d\alpha(t)}{dt}v(\alpha)$, which is the same form as Eq. (7). It has also been verified that an optimal functional form of $\alpha(t)$ is obtained by minimizing the maximum value of the inverse traceability grover_optimal , which ensures a consistent traceability throughout the interpolation process. In CRS, we regard the probability-distributional change as the inverse traceability and minimize the maximum distance for determining the noise schedule.
4.3 Distance Metrics
Any distance metric is applicable to CRS. We introduce three distance metrics.
Distance metric based on FID: FID is a widely used distance metric to evaluate the sample fidelity in the field of image generation. The FID calculates the Fréchet distance fid_score under a normality assumption in the vision-relevant feature space embedded using the Inception-V3 model inception_v3 , which is trained on the ImageNet classification. When FID is used as the distance metric for the pixel-space diffusion models, CRS allocates more timesteps to the region where the sample fidelity significantly deteriorates in the diffusion process ($\alpha\simeq 1$). A weakness of FID-based distance metric is that it cannot accurately capture the probability-distributional change in the region with high-noise levels ($\alpha\simeq 0$), because the feature model is not trained on diffused images.
It is important to note that, in the latent-space diffusion models, the probability-distributional change must be measured in the latent space embedded using the autoencoder (the approximation in Eq. (3) is applied to the diffusion process defined in the latent space). We trained ResNet50 resnet on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model for computing FID.
We refer to $v(\alpha)$ computed using FID as $v_{\mathrm{FID}}(\alpha)$. Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain the noise schedule using Eq. (8). For the detailed computation procedure of $v_{\mathrm{FID}}(\alpha)$, refer to Appendix C.1.
Distance metric based on data prediction: We introduce a distance metric that uses a trained diffusion model as follows:
$v_{x}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{x}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].\$ (9)
Here, $\bm{x}_{\alpha}$ and $\bm{x}_{\alpha^{\prime}}$ are sampled from $q(\bm{x}_{\alpha},\bm{x}_{\alpha}^{\prime})=q(\bm{x}_{\alpha}^{\prime}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})$, and $\bm{x}_{\theta}(\bm{x},\alpha)$ is the data prediction, which is related to the noise prediction $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ as follows: $\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$. This distance metric aims at allocating more computational resources to the region where the data prediction changes rapidly. As shown in Appendix B, we can interpret this distance metric as a simplified version of the Kullback-Leibler divergence upper bound (KLUB) introduced in AYS. This distance metric can be used not only for determining the sampling schedule after training a diffusion model, but also for adaptively optimizing the training schedule during the training process. Pseudocodes for optimizing training and sampling schedules using $v_{x}(\alpha)$ are shown in Appendix C.2.
We can interpret that $v_{x}(\alpha)$ measures the probability-distributional change of diffused data. The data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ is given by the weighted sum of the data points $\bm{x}_{0}$ statistical_mechanics . The weight of each data point depends on the noise level $\alpha$. The weights of all data points are uniform when $\alpha=0$, while a small number of data points contribute to $\bm{x}_{\theta}(\bm{x},\alpha)$ when $\alpha\simeq 1$. The changes in these weights represent how the trajectories in the diffusion process gradually diverge as the noise level decreases. Therefore, we can interpret $v_{x}(\alpha)$ as measuring the probability-distributional change of diffused data.
Unlike the FID-based distance metric, $v_{x}(\alpha)$ can capture distributional changes even in regions with high-noise levels. However, $v_{x}(\alpha)$, which is related to the KLUB, is not necessarily optimal for improving sample fidelity at sampling. Therefore, as discussed in the next subsection, combining multiple distance metrics in a complementary manner enhances the performance of CRS.
Distance metric based on noise prediction: In the experiments presented in Sec. 5, we trained the noise prediction model. In this case, it is also considered effective to use noise predictions to measure the probability-distributional change as follows:
$v_{\varepsilon}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (10)
Hereafter, we refer to CRS using $v_{\mathrm{FID}}(\alpha)$, $v_{x}(\alpha)$, and $v_{\varepsilon}(\alpha)$ for optimizing noise schedules as CRS-$v_{\mathrm{FID}}$, CRS-$v_{x}$, and CRS-$v_{\varepsilon}$, respectively.
4.4 Combining Multiple Distance Metrics
As explained in the previous subsection, each distance metric has its own strengths an weaknesses. Therefore, it is considered effective to complementarily combining multiple distance metrics for optimizing noise schedules. We introduce a strategy for combining them as follows:
$v(\alpha)=\sum_{m=1}^{M}w_{m}\tilde{v}_{m}(\alpha),\ \ \mathrm{where}\ \ \tilde{v}_{m}(\alpha)=\frac{1}{C_{m}}v_{m}(\alpha)^{\xi_{m}}\ \ \mathrm{and}\ \ C_{m}=\int_{\alpha_{\min}}^{\alpha_{\max}}v_{m}(\alpha)^{\xi_{m}}.$ (11)
Here, $M$ is the number of distance metrics to be combined, $v_{m}(\alpha)$ is computed for each distance metric, $w_{m}$ is the weight for $v_{m}(\alpha)$ satisfying $\sum_{m=1}^{M}w_{m}=1$, and $\xi_{m}$ controls the dependence on $v_{m}(\alpha)$. Note that the value range of $v_{m}(\alpha)$ can significantly vary depending on the distance metric. Therefore, normalization of $v_{m}(\alpha)$ is required before combining them, as shown in Eq. (11).
Using this formulation, it is also possible to combine conventional noise schedules with CRS based noise schedules by converting $\alpha(t)$ to $v(\alpha)$. For example, when $\alpha(t)=\cos\left(\frac{\pi t}{2}\right)$, $v(\alpha)$ is obtained as follows:
$v(\alpha)\propto-\left(\frac{d\alpha(t)}{dt}\right)^{-1}=\frac{2}{\pi}\sin^{-1}\left(\frac{\pi t}{2}\right)=\frac{2}{\pi\sqrt{1-\alpha^{2}}}.$ (12)
This result suggests that every noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
5 Experiments
We experimentally demonstrated that CRS broadly improves the performance of diffusion models. Below, we provide a brief explanation of each component of the diffusion model.
Dataset: We use six image datasets: LSUN (Church, Bedroom, Horse) dataset_lsun , ImageNet dataset_imagenet , FFHQ dataset_ffhq , and CIFAR10 dataset_cifar10 . LSUN and FFHQ are used for unconditional image generation, and ImageNet and CIFAR10 for class-conditional image generation.
Noise Prediction Model: We use the U-Net architecture introduced in the ablated diffusion model (ADM) adm as the noise prediction model (see Table 8 in Appendix E for hyperparameter settings).
Training Schedule: We use linear ddpm_original , shifted cosine shifted_cosine , adaptive noise schedule proposed in VDM++ vdmpp , and CRS.
Loss Function: Except for the adaptive noise schedule of VDM++, we use the following simplified loss function introduced in DDPMs:
$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},t\sim\mathcal{U}(0,1),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x},\alpha(t))-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (13)
where $\mathcal{D}$ denotes all data points in the training dataset, $\alpha(t)$ is the training schedule, and $\bm{x}=\alpha(t)\bm{x}_{0}+\sqrt{1-\alpha(t)^{2}}\bm{\varepsilon}$.
When using the adaptive noise schedule of VDM++, the following loss function is applied:
$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\lambda\sim p(\lambda),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\frac{\omega(\lambda)}{p(\lambda)}\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (14)
where $\lambda=\log\left(\frac{\alpha^{2}}{\sigma^{2}}\right)$ is the log signal-to-noise ratio (log-SNR), $\omega(\lambda)$ is the weight for each noise level, and $p(\lambda)=-\frac{d\lambda}{dt}$ is the noise schedule for training. We use the EDM monotonic function as $\omega(\lambda)$, and $p(\lambda)$ is updated at each iteration to approximately satisfy the following condition:
$p(\lambda)\propto\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\omega(\lambda)\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right].$ (15)
Sampling Schedule: We use linear, shifted cosine, EDM sampling edm and CRS. In pixel-space diffusion models, we additionally evaluate CRS-$v_{x}+v_{\mathrm{FID}}$, which combines $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ for optimizing noise schedules. The reason for selecting the combination of $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ is that, as shown in Sec. 5.4, $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ have a complementary relationships. Recent studies edm ; vdmpp have shown that using different noise schedules for training and sampling can lead to improved performance. Therefore, in our experiments, we evaluate the performance of the diffusion models using arbitrary pairs of training and sampling schedules. Note that we can obtain different noise schedules for training and sampling using CRS, because it is capable of handling arbitrary distance metrics. Details of the conventional noise schedules used in our experiments can be found in Table 10 in Appendix E.
Sampling Methods: Images are mainly generated using two stochastic samplers and three deterministic samplers. For the stochastic samplers, we use Stochastic DDIM ($\eta=1$) ddim_original and SDE-DPM-Solver++(2M) sampler_dpmpp . For the deterministic samplers, we use DDIM ddim_original , DPM-Solver++(2M) sampler_dpmpp , and pseudo numerical methods for diffusion models (PNDMs) sampler_pndm . At small NFEs, we additionally use improved PNDM (iPNDM) sampler_ipndm and a unified predictor-corrector framework (UniPC) sampler_unipc .
Evaluation Metrics: We compute four metrics: FID fid_score , spatial FID (sFID) spatial_fid_score , and improved precision and recall recall_precision1 ; recall_precision2 . FID, sFID, and precision measure sample fidelity, while recall evaluates mode coverage.
5.1 Results on Pixel-Space Diffusion Models
We trained the continuous-time diffusion models on LSUH Horse 256 $\times$ 256 and Bedroom 256 $\times$ 256 for unconditional image generation. Detailed experimental results are listed in Appendix F.1.
LSUN Horse 256 $\times$ 256: The evaluation results are summarized in Table 2. The rows shaded in gray represent the dependence on the training schedule. Although sample fidelity and mode coverage are generally considered to be in a trade-off relationship, CRS-$v_{x}$ achieved the highest performance for all metrics. This result indicates that CRS-$v_{x}$ is not merely balancing trade-offs between evaluation metrics but fundamentally improves the quality of the generated samples.
The unshaded rows represent the dependence on sampling schedules. CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the highest performance for all metrics, indicating fundamental improvement in sample quality (see Appendix D for the hyperparameter setting of CRS-$v_{x}+v_{\mathrm{FID}}$). The FID score of 2.03 surpasses the current state-of-the-art FID score of 2.11 lsun_horse_sota .
LSUN Bedroom 256 $\times$ 256: We confirmed the advantage of CRS in the low-NFE regime by comparing it with the leading results reported in GITS noise_gits . We trained the continuous-time diffusion model and used CRS-$v_{x}$ for optimizing the training schedule. The results are listed in Table 2. The best FID scores at $\mathrm{NFE}=5$ and $\mathrm{NFE}=10$ were achieved when using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$ for optimizing the training and sampling schedules, respectively.
Table 1: Dependence on training and sampling schedules for LSUN Horse $256\times 256$ in pixel-space diffusion models at $\mathrm{NFE}=250$. SDE-DPM-Solver++(2M) was used as sampler. Bold values represent best training or sampling schedules for each metric.
Training Sampling Metrics
schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Linear Linear 2.90 6.82 0.66 0.56
Shifted cosine 2.95 6.43 0.67 0.56
VDM++ 3.82 7.40 0.66 0.54
CRS-$v_{\varepsilon}$ 2.91 6.56 0.66 0.56
CRS-$v_{x}$ 2.73 6.40 0.67 0.56
CRS-$v_{x}$ Linear 2.73 6.40 0.67 0.56
Shifted cosine 3.13 7.09 0.67 0.54
EDM 2.09 6.16 0.69 0.56
CRS-$v_{\varepsilon}$ 2.87 6.71 0.66 0.56
CRS-$v_{x}$ 5.46 8.34 0.62 0.52
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.03 6.06 0.69 0.56
Table 2: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion models at small NFEs. Bold values indicate best FID scores at each NFE. Results for rows other than "Ours" are taken from previously published study noise_gits .
Model Sampler Sampling schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$
EDM UniPC sampler_unipc EDM 23.34 5.75
AMED-Solver sampler_amed 5.38
iPNDM sampler_ipndm GITS noise_gits 15.85 5.28
Ours DPM-Solver++(2M) EDM 23.72 4.73
CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 4.88
UniPC EDM 82.52 4.94
CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 3.30
5.2 Results on Latent-Space Diffusion Models
We trained latent-space diffusion models using the pre-trained autoencoder VQ-4 used in latent diffusion models (LDMs) ldm . We used LSUN Church and Bedroom for unconditional image generation, and ImageNet for class-conditional image generation. The image resolution was set to $256\times 256$, which was reduced to $64\times 64$ in the latent space. We present the evaluation results on LSUN Church $256\times 256$. Detailed results are listed in Appendix F.2.
LSUN Church 256 $\times$ 256: The evaluation results are summarized in Table 4. The dependence on the training schedule is shown in the rows shaded in gray, and CRS-$v_{x}$ achieved the highest performance except for sFID. The unshaded rows represent the dependence on the sampling schedule, and CRS-$v_{x}$ further improves all evaluation metrics except for precision. These results indicate the superiority of CRS in latent-space diffusion models as well.
From the results in Table 4, CRS-$v_{x}$ appears to perform the best at both training and sampling; however, depending on the sampler and NFE, CRS-$v_{\varepsilon}$ is preferable at sampling.
5.3 Comparison with AYS Using EDM
Because $v_{x}(\alpha)$ is related to the KLUB introduced in AYS, we additionally compare CRS with AYS on FFHQ 64 $\times$ 64 using the pretrained model provided in EDMs edm . It is important to note that EDM adopts a variance-exploding process, while we have so far focused on the variance-preserving process. However, CRS can also be applied to the variance-exploding process by converting between $\alpha$ and the noise level in EDM $\sigma_{\mathrm{EDM}}$ via the log SNR, as shown below:
$\mathrm{logSNR}=\log\frac{\alpha^{2}}{1-\alpha^{2}}=\log\frac{1}{\sigma_{\mathrm{EDM}}^{2}}\Rightarrow\alpha=\frac{1}{\sqrt{1+\sigma_{\mathrm{EDM}}^{2}}}.$ (16)
The results are listed in Table 4. CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperformed AYS at all NFEs. Detailed results on pretrained EDM are listed in Appendix F.3.
Table 3: Dependence on training and sampling schedules for LSUN Church $256\times 256$ in latent-space diffusion models at $\mathrm{NFE}=30$. DPM-Solver++(2M) was used as sampler. Bold values represent best training or sampling schedules for each metric.
Training Sampling Metrics
schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Linear Linear 3.82 10.28 0.60 0.58
VDM++ 4.00 10.07 0.58 0.59
CRS-$v_{\mathrm{FID}}$ 3.78 10.51 0.60 0.58
CRS-$v_{\varepsilon}$ 3.87 10.30 0.60 0.59
CRS-$v_{x}$ 3.69 10.23 0.60 0.59
CRS-$v_{x}$ Linear 3.69 10.23 0.60 0.59
EDM 4.31 11.31 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.98 9.61 0.58 0.60
CRS-$v_{\varepsilon}$ 3.63 9.91 0.59 0.60
CRS-$v_{x}$ 3.59 9.51 0.59 0.60
Table 4: FID scores evaluated on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM. FID scores on AYS is taken from their original paper noise_ays .
Sampler Noise Schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) AYS noise_ays 5.43 3.29 2.62
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
5.4 Comparison of Noise Schedules
The CRS-based noise schedules in the pixel-space diffusion model are shown in Fig. 3. The hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$ was set to $w_{x}=w_{\mathrm{FID}}=0.5$ and $\xi_{x}=\xi_{\mathrm{FID}}=1.0$. Since $v_{\mathrm{FID}}(\alpha)$ takes large values in regions where sample fidelity degrades in the diffusion process, CRS-$v_{\mathrm{FID}}$ allocates many timesteps to the region where $\alpha\simeq 1$. In contrast, CRS-$v_{x}$ allocates more timesteps to regions with small $\alpha$. This is due to the fact that the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ can be expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$ using the noise prediction, where $\alpha$ appears in the denominator. Thus, CRS-$v_{\mathrm{FID}}$ and CRS-$v_{x}$ exhibits complementary characteristics, and combining them yields an effective sampling schedule. Since the noise schedule based on CRS-$v_{\varepsilon}$ can be approximately reproduced by tuning the hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$, we regard $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ as the fundamental compenents.
Figure 3 shows the CRS-based noise schedules in the latent-space diffusion model. The appropriate noise schedules differ significantly between the pixel-space and lantet-space diffusion models. The most notable difference is that, in the latent-space diffusion models, fewer timesteps are allocated to the region where $\alpha\simeq 1$. This can be attributed to the fact that, in latent-space diffusion models, the perceptual expression is largely generated by the autoencoder, and the autoencoder can absorb minor deviations in the generated latent variables. In contrast, the pixel-space diffusion models directly generate images, so careful control of the noise schedule is crucial at $\alpha\simeq 1$. We did not evaluate any combination of multiple distance metrics in latent-space diffusion models, because all the CRS-based noise schedules have a similar functional form.
Figure 2: Noise schedules on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model.
Figure 3: Noise schedules on LSUN Church 256 $\times$ 256 in latent-space diffusion model.
6 Limitations
This paper focuses on developing CRS and validating its effectiveness under fundamental experimental settings. However, to further strengthen the generality of CRS, additional investigations are needed from the following perspectives: 1) generality across various generation conditions, 2) scalability of the distance metrics with respect to dataset size and data dimensionality, 3) effectiveness in domains beyond image generation. While $v_{\mathrm{FID}}(\alpha)$ proved effective as the sampling schedule in pixel-space diffusion models, it is a distance metric specialized for image generation. Therefore, when applying CRS to other domains, it may be necessary to introduce new distance metrics that are better suited to the characteristics of the target domain.
7 Conclusion
We proposed a general approach called CRS to optimize noise schedules for training and sampling in diffusion models. Through extensive experiments, we confirmed that CRS broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and NFEs ranging from 5 to 250. CRS is fundamentally different from conventional approaches noise_ays ; noise_gits in that it can optimize the training schedule as well and use any distance metric for measuring probability-distributional change, allowing the introduction of domain specific distance metric and the use of different noise schedules for training and sampling. Although this study has the limitations discussed in Sec. 6, we believe that it successfully provides solid baselines. We hope that future research will further improve the generality of CRS, leading to its widespread adoption as a general-purpose approach.
Appendix A Probability Distribution of Diffused Data
We derive the probability distribution of diffused data used in the toy example (Fig. 1).
The probability distribution of diffused data is given by
[…]
$q(\bm{x}_{\alpha}|\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0},\sigma^{2}\bm{I}),$ (18)
where $q(\bm{x}_{0})$ is the probability distribution of a target dataset. We then approximate $q(\bm{x}_{0})$ with the empirical distribution as
$q(\bm{x}_{0})=\frac{1}{N}\sum_{n=1}^{N}\delta(\bm{x}_{0}-\bm{x}_{0}^{n}),$ (19)
where $N$ is the number of samples in the target dataset, $\bm{x}_{0}^{n}$ is the $n$-th sample, and $\delta$ denotes Dirac’s delta function. By substituting Eq. (19) into Eq. (17), we obtain
$q(\bm{x}_{\alpha})=\frac{1}{N}\sum_{n=1}^{N}\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0}^{n},\sigma^{2}\bm{I}).$ (20)
We used the following formula of Dirac’s delta function:
[…]
The resulting probability distribution is a Gaussian mixture distribution. We used Eq. (20) in the toy example.
Appendix B Derivation of Distance Metric Based on Data Prediction from KLUB
We show that $v_{x}(\alpha)$ can be derived as a simplified version of KLUB introduced in AYS.
B.1 KLUB in Variance-Preserving Process
The forward SDE of the continuous-time diffusion models [19] is defined as
$d\bm{x}=f(t)\bm{x}dt+g(t)d\bm{\omega},$ (22)
where $f(t)\bm{x}$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $\bm{\omega}$ is the standard Wiener process. The associated reverse SDE is given by
$d\bm{x}=\left[f(t)\bm{x}-g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\right]dt+g(t)d\bm{\omega}.$ (23)
For approximately computing the score $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, we train the following noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$:
$\bm{\varepsilon}_{\theta}(\bm{x},\alpha)\simeq-\sigma\nabla_{\bm{x}}\log p_{t}(\bm{x}).$ (24)
The noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ has the following relationship with the data prediction model $\bm{x}_{\theta}(\bm{x},\alpha)$:
$\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha).$ (25)
By substituting Eq. (24) into Eq. (23) and using Eq. (25), the reverse SDE can be rewritten as
$d\bm{x}$ $=\bm{f}(\bm{x},t)dt+g(t)d\bm{\omega},$ (26)
$\bm{f}(\bm{x},t)$ $=\left(f(t)+\frac{1}{\sigma(t)^{2}}g(t)^{2}\right)\bm{x}-\frac{\alpha(t)}{\sigma(t)^{2}}g(t)^{2}\bm{x}_{\theta}(\bm{x},\alpha(t)).$ (27)
Stochastic DDIM is a stochastic sampler designed to numerically solve the reverse SDE described above. In Stochastic DDIM, the solution for the first term of $\bm{f}(\bm{x},t)$ (linear term) is exactly computed, and discretization errors arise only from the second term of $\bm{f}(\bm{x},t)$ [22, 23]. According to a previous study [16], in the sub-interval $[t_{i-1},t_{i}]$, Stochastic DDIM solves the following discretized SDE:
$d\bm{x}$ $=\bm{f}^{\prime}(\bm{x},t)dt+g(t)d\bm{\omega},$ (28)
$\bm{f}^{\prime}(\bm{x},t)$ $=\left(f(t)+\frac{1}{\sigma(t)^{2}}g(t)^{2}\right)\bm{x}-\frac{\alpha(t)}{\sigma(t)^{2}}g(t)^{2}\bm{x}_{\theta}(\bm{x}_{t_{i}},\alpha(t_{i})).$ (29)
The KLUB between probability distributions of $\bm{x}_{t_{i-1}}$ obtained by exactly solving the reverse SDE and approximately solving the reverse SDE using Stochastic DDIM is introduced as
[…]
The diffusion coefficient $g(t)$ in the variance-preserving process is given by
$\frac{1}{2}g(t)^{2}=-\frac{\dot{\alpha}}{\alpha}.$ (32)
By substituting Eq. (32) into Eq. (31), we obtain
[…]
B.2 KLUB as Distance Metric of CRS
We transform the KLUB into a form that can be applied to CRS. Because the KLUB is used to evaluate the discretization error when solving the reverse SDE, it is reasonable to regard the KLUB as the traceability measure of the diffusion process.
First, we approximate the integral in the KLUB using Simpson’s rule as follows:
$\mathrm{KLUB}(t,t+\Delta t)$ $=\frac{\dot{\alpha(t)}\alpha(t)}{2\sigma(t)^{4}}\overline{\mathcal{D}_{x}^{2}}(t,t+\Delta t)\Delta t,$ (34)
$\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ $=\mathbb{E}_{\begin{subarray}{c}\bm{x}_{0}\sim q(\bm{x}_{0})\\ \bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})\\ \bm{x}_{t^{\prime}}\sim q(\bm{x}_{t^{\prime}}|\bm{x}_{t})\end{subarray}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{t},\alpha(t))-\bm{x}(\bm{x}_{t^{\prime}},\alpha(t^{\prime}))\right\|_{2}^{2}\right].$ (35)
Next, to treat $\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$ as a function of $\alpha$ rather than $t$, we introduce $\overline{D_{x}^{2}}(\alpha(t),\alpha(t^{\prime}))=\overline{\mathcal{D}_{x}^{2}}(t,t^{\prime})$:
$\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (36)
Here, $q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ is given by
$q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ $=q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0}),$ (37)
$q(\bm{x}_{\alpha}|\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{\alpha};\alpha\bm{x}_{0},\sigma^{2}\bm{I}),$ (38)
$q(\bm{x}_{\alpha^{\prime}}|\bm{x}_{\alpha})$ $=\mathcal{N}(\bm{x}_{\alpha^{\prime}};\beta\bm{x}_{\alpha},\delta^{2}\bm{I}),$ (39)
where $\beta=\frac{\alpha^{\prime}}{\alpha}$, $\sigma=\sqrt{1-\alpha^{2}}$, and $\delta=\sqrt{1-\beta^{2}}$. Using $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$, we can rewrite $\mathrm{KLUB}(t,t+\Delta t)$ as
$\mathrm{KLUB}(t,t+\Delta t)$ $=\frac{\dot{\alpha}\alpha}{2\sigma^{4}}\overline{D_{x}^{2}}(\alpha,\alpha+\Delta\alpha)\Delta t,$ (40)
$=\frac{\dot{\alpha}\alpha}{2\sigma^{4}}\frac{\overline{D_{x}^{2}}(\alpha,\alpha+\Delta\alpha)-\overline{D_{x}^{2}}(\alpha,\alpha)}{\Delta\alpha}\frac{\Delta\alpha}{\Delta t}(\Delta t)^{2},$ (41)
$\simeq\left\{\frac{d\alpha}{dt}\cdot\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}\sqrt{\left.\frac{\partial\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\Delta t\right\}^{2}.$ (42)
We use $\overline{D_{x}^{2}}(\alpha,\alpha)=0$ in the second equality. Finally, by adopting $\sqrt{\mathrm{KLUB}(t,t^{\prime})}$ as the distance metric $D(t,t^{\prime})$ in CRS, we obtain the following equation:
$D(t,t+\Delta t)=\sqrt{\mathrm{KLUB}(t,t+\Delta t)}=-\frac{d\alpha}{dt}\cdot\underbrace{\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}\sqrt{\left.\frac{\partial\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}}_{v_{\mathrm{KLUB}}(\alpha)}\Delta t.$ (43)
which is the same form as Eq. (7). We can regard $v_{x}(\alpha)$ as a simplified version of $v_{\mathrm{KLUB}}(\alpha)$ in the sense that the weighting function of $v_{\mathrm{KLUB}}(\alpha)$, namely $\frac{\sqrt{\alpha}}{\sqrt{2}\sigma^{2}}$, is discarded in $v_{x}(\alpha)$.
It is possible to optimize the noise schedule using $v_{\mathrm{KLUB}}(\alpha)$. However, when $v_{\mathrm{KLUB}}(\alpha)$ is directly used, the noise schedule is heavily impacted by the weighting function rather than $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$. The weighting function becomes extremely large around $\alpha\simeq 1(\sigma\simeq 0)$, which overwhelms the impact of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$. Recall that $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ can be interpreted as measuring probability-distributional changes of diffused data, as described in Sec. 4.3. Therefore, the simplified version we introduce is more aligned with our objective of optimizing the noise schedule on the basis of distributional changes.
Appendix C Implementation of CRS
We describe how to compute $v(\alpha)$ using each distance metric.
C.1 Distance Metric Based on FID
The forward process is simulated using the training dataset for computing $v_{\mathrm{FID}}(\alpha)$, as shown in Algorithm 1. One of the inputs $\{\alpha_{t}\}_{t=0,...,T}$ needs to be discretized to approximately represent the functional form of $v(\alpha)$ through linear interpolation. In our experiments, we set
$\alpha_{t}=1-\left(\frac{t}{T}\right)^{p},$ (44)
where $T=1000$, and set $p$ to $2$ and $1$ for the diffusion processes in pixel space and latent space, respectively. For the feature model $\bm{\phi}$, we used the Inception-V3 model trained on ImageNet classification for the pixel-space diffusion process. For the latent-space diffusion process, however, we need a feature model trained in the latent space embedded using the autoencoder. We trained ResNet50 on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model $\bm{\phi}$. Note that a trained diffusion model is not necessary for computing $v_{\mathrm{FID}}(\alpha)$. Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain noise schedules by solving Eq. (8).
The dependence of CRS-$v_{\mathrm{FID}}$ on hyperparameter $\xi$ is illustrated in Fig. 7. CRS-$v_{\mathrm{FID}}$ is computed on LSUN Horse 256 $\times$ 256 for the pixel-space diffusion process. As $\xi$ is increased, more timesteps are allocated to the region with low-noise level.
The computation time of $v_{\mathrm{FID}}(\alpha)$ is proportional to the number of images in the training dataset. For the pixel-space diffusion process, it takes about 35 hours for LSUN Horse 256 $\times$ 256 with 2M images using eight A100 GPUs. For the latent-space diffusion process, it takes about 12 hours for ImageNet 256 $\times$ 256 with 1.28M images using eight A100 GPUs. The computation of $v_{\mathrm{FID}}(\alpha)$ needs to be executed only once before training the diffusion model, and the computation time is sufficiently small compared with the training time of the diffusion model.
Algorithm 1 $v(\alpha)$ on basis of FID
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, number of timesteps for simulating the forward process $T$, discretized noise levels $\{\alpha_{t}\}_{t=0,...,T}$, feature model $\bm{\phi}$
[…]
1:  # Compute statistics of feature vectors
2:  for $n=0$ to $N-1$ do
3:     $\bm{x}^{(n)}_{0}\leftarrow\bm{x}^{(n)}$
4:     if diffusion process in latent space then
5:        $\bm{x}^{(n)}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}^{(n)}_{0})$
6:     end if
7:
8:     # Simulate the forward process
9:     for $t=1$ to $T$ do
10:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}}$
11:        $\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
12:        $x^{(n)}_{t}\leftarrow\beta_{t}\bm{x}^{(n)}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
13:     end for
14:  end for
15:  $\bm{\mu}_{t}\leftarrow mean\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
16:  $\bm{\Sigma}_{t}\leftarrow cov\left(\{\bm{x}^{(n)}_{t}\}_{n=0,...,N-1}\right),\ \forall{t}\in\{0,...,T\}$
17:
18:  # Compute $v(\alpha)$ using $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$
19:  for $t=0$ to $T-1$ do
20:     $v_{t}\leftarrow\frac{\mathrm{FID}(\bm{\mu}_{t},\bm{\Sigma}_{t},\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})}{\alpha_{t}-\alpha_{t+1}}$
21:  end for
22:  $v_{T}\leftarrow v_{T-1}$
Figure 4: Dependence of CRS-$v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$.
Figure 5: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $w_{x}$ and $w_{\mathrm{FID}}$ ($\xi_{x}=\xi_{\mathrm{FID}}=1)$.
Figure 6: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{\mathrm{FID}}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
Figure 7: Dependence of CRS-$v_{x}+v_{\mathrm{FID}}$ on $\xi_{x}$ ($w_{x}=w_{\mathrm{FID}}=0.5)$.
C.2 Distance Metric Based on Data Prediction
Algorithm 2 presents the pseudocode for computing $v_{x}(\alpha)$ using a trained diffusion model. This pseudocode focuses on the optimization of sampling schedules. Similar to the computation of $v_{\mathrm{FID}}(\alpha)$, the forward process is simulated using the training dataset, and the change in the data prediction between adjacent timesteps is computed. A common configuration of hyperparameters $(T=10^{3},S=10^{4},\alpha_{s}=1,\alpha_{e}=10^{-4})$ was used in our experiments.
The computation time for obtaining $v_{x}(\alpha)$ is negligible compared with the training time of the diffusion model. It takes approximately 5.5 hours using eight H100 GPUs in pixel-space diffusion models. In latent-space diffusion models, it takes approximately 1.5 hours using eight A100 GPUs.
Algorithm 2 $v_{x}(\alpha)$ for optimizing sampling schedule.
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, trained diffusion model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$, number of timesteps for simulating the forward process $T$, number of samples used for calculating the mean $S$, maximum and minimum values of $\alpha$ $(\alpha_{s},\alpha_{e})$
0:  $LinearInterp1d(\{\alpha_{t}\}_{t=0,...,T},\{v_{t}\}_{t=0,...,T})$
1:  # Compute $\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right]$
2:  $\overline{D^{2}_{t}}\leftarrow 0,\ \forall{t}\in\{0,1,...,T\}$
3:  $\Delta\alpha\leftarrow\frac{1}{T}(\alpha_{s}-\alpha_{e})$
4:  for $n$ in $randperm(N)[:S]$ do
5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)}$
6:     if latent-space diffusion model then
7:        $\bm{x}_{0}\leftarrow autoencoder(\bm{x}_{0})$
8:     end if
9:     $\bm{y}_{0}\leftarrow\bm{x}_{0}$
10:
11:     # Simulate the forward process
12:     for $t=1$ to $T$ do
13:        $\alpha_{t}\leftarrow\alpha_{s}-\Delta\alpha t,\ \sigma_{t}\leftarrow\sqrt{1-\alpha_{t}^{2}}$
14:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}},\ \delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
15:        $\bm{x}_{t}\leftarrow\beta_{t}\bm{x}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
16:        $\bm{y}_{t}\leftarrow\frac{1}{\alpha_{t}}\left(\bm{x}_{t}-\sigma_{t}\bm{\varepsilon}_{\theta}(\bm{x}_{t},\alpha_{t})\right)$
17:        $\overline{D^{2}_{t}}\leftarrow\overline{D^{2}_{t}}+\frac{1}{S}\left\|\bm{y}_{t}-\bm{y}_{t-1}\right\|_{2}^{2}$
18:     end for
19:  end for
20:
21:  # Compute $v(\alpha)$ using $\overline{D^{2}_{t}}$
22:  for $t=1$ to $T$ do
23:     $v_{t}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{t}}}$
24:  end for
25:  $v_{0}\leftarrow v_{1}$
CRS-$v_{x}$ can be used for optimizing the training schedule as well. A pseudocode for training diffusion models with CRS is given in Algorithm 3. Since the data prediction is expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, $v_{x}(\alpha)$ tends to take extremely large values when $\alpha\simeq 0$. To avoid this, $v_{x}(\alpha)$ is evaluated within the range of $\alpha_{\mathrm{th}}\leq\alpha\leq\alpha_{\max}$, and we set $v_{x}(\alpha<\alpha_{\mathrm{th}})=v_{x}(\alpha_{\mathrm{th}})$. Similar to the implementation of the adaptive noise schedule in VDM++ [13], the range of $\alpha$ from $\alpha_{\mathrm{th}}$ to $\alpha_{\max}$ is divided into $B$ evenly spaced bins, and $\overline{D^{2}_{x}}(\alpha,\alpha+\Delta\alpha)$ is approximately evaluated using the exponential moving average (EMA). We used a common configuration of hyperparameters ($B=100,e=0.995,\alpha_{\max}=1,\alpha_{\min}=0,\alpha_{\mathrm{th}}=0.01,\Delta\alpha=10^{-3},\xi=1.0)$ in this study. The initial noise schedule is set to $\alpha(t)=\alpha_{\max}-(\alpha_{\max}-\alpha_{\min})\frac{t}{T}$. Although not explicitly included in the pseudocode, the update of $\overline{D^{2}_{b}}$ starts after 1000 iterations, and the noise schedule $\alpha(t)$ is updated every 100 iterations.
The training times of the diffusion models are listed in Table 5. When using CRS for optimizing training schedule, the training time increases by 20–30% compared with conventional noise schedules because CRS requires evaluating $v_{x}(\alpha)$ during training. To mitigate the increase in training time, one potential approach is to fix the training schedule after a certain number of epochs. In practice, CRS-$v_{x}$ showed little change beyond 10 epochs on LSUN Horse 256 $\times$ 256, suggesting the possibility of reducing training time while maintaining performance by fixing the training schedule after 10 epochs. Reducing the update frequency of $\overline{D^{2}_{x}}$ can be also an effective strategy. Developing an efficient implementation to reduce the training time remains as future work.
Algorithm 3 Adaptive noise schedule for training.
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N}$, initial noise schedule $\alpha(t)$, number of bins $B$, decay rate of EMA $e$, maximum and minimum values of $\alpha$ for training schedule $(\alpha_{\max},\alpha_{\min})$, infinitesimal variation of $\alpha$ for evaluating the derivative $\Delta\alpha$, minimum $\alpha$ for evaluating $v(\alpha)$ $\alpha_{\mathrm{th}}$, hyperparameter of CRS $\xi$
0:  noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$
1:  $\alpha_{b}\leftarrow\alpha_{\mathrm{th}}+(\alpha_{\max}-\alpha_{\mathrm{th}})\frac{b}{B},\ \forall{b}\in\{0,...,B\}$
2:  $\overline{D^{2}_{b}}\leftarrow 1e^{-6},\ \forall{b}\in\{0,...,B-1\}$
3:  for $n$ in $randperm(N)$ do
4:     # Compute loss for training diffusion models
5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)},\ t\sim\mathcal{U}(0,1),\ \bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
6:     $\alpha\leftarrow\alpha(t),\ \sigma\leftarrow\sqrt{1-\alpha^{2}}$
7:     $\bm{\varepsilon}_{\alpha}\leftarrow\bm{\varepsilon}_{\theta}(\alpha\bm{x}_{0}+\sigma\bm{\varepsilon},\alpha)$
8:     $loss\leftarrow\frac{1}{2}\left\|\bm{\varepsilon}_{\alpha}-\bm{\varepsilon}\right\|_{2}^{2}$
9:
10:     # Evaluate $\overline{D^{2}_{x}}$ using EMA (without gradient)
11:     if $\alpha\geq\alpha_{\mathrm{th}}$ then
12:        $\alpha^{\prime}\leftarrow\alpha-\Delta\alpha,\ \beta^{\prime}\leftarrow\frac{\alpha^{\prime}}{\alpha},\ \delta^{\prime}\leftarrow\sqrt{1-\beta^{\prime 2}}$
13:        $\bm{x}_{\alpha^{\prime}}\leftarrow\beta^{\prime}\bm{x}_{\alpha}+\delta^{\prime}\times\mathcal{N}(\bm{0},\bm{I})$
14:        $\bm{\varepsilon}_{\alpha^{\prime}}\leftarrow\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})$
15:        $\bm{y}_{\alpha}\leftarrow\frac{1}{\alpha}\left(\bm{x}_{\alpha}-\sigma\bm{\varepsilon}_{\alpha}\right),\ \bm{y}_{\alpha^{\prime}}\leftarrow\frac{1}{\alpha^{\prime}}\left(\bm{x}_{\alpha^{\prime}}-\sigma^{\prime}\bm{\varepsilon}_{\alpha^{\prime}}\right)$
16:        $b\leftarrow\min\left(\left\lfloor\frac{\alpha-\alpha_{\mathrm{th}}}{\alpha_{\max}-\alpha_{\mathrm{th}}}B\right\rfloor,B-1\right)$
17:        $\overline{D^{2}_{b}}\leftarrow e\overline{D^{2}_{b}}+(1-e)\left\|\bm{y}_{\alpha}-\bm{y}_{\alpha^{\prime}}\right\|_{2}^{2}$
18:        $v_{b}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{b}}}$
19:     end if
20:
21:     # Update training schedule $\alpha(t)$
22:     $\mathcal{X}\leftarrow\{\alpha_{\min},\alpha_{0},\alpha_{1},...,\alpha_{B}\}$
23:     $\mathcal{Y}\leftarrow\{v_{0},v_{0},v_{1},...,v_{B-1},v_{B-1}\}$
24:     $v(\alpha)\leftarrow LinearInterp1d(\mathcal{X},\mathcal{Y})$
25:     Update $\alpha(t)$ and model parameters $\theta$
26:  end for
Table 5: Training time of diffusion models.
Diffusion type Dataset GPUs Epochs Training schedule Training time
Latent space LSUN Church A100 $\times$ 8 1000
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
85 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
105 hours
LSUN Bedroom A100 $\times$ 8 100
CRS-$v_{\mathrm{FID}}$
190 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
245 hours
ImageNet A100 $\times$ 8 500
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
430 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
525 hours
Pixel space LSUN Horse H100 $\times$ 8 50
Linear
Shifted cosine
VDM++
170 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
215 hours
LSUN Bedroom H100 $\times$ 8 50
CRS-$v_{x}$
325 hours
Appendix D Hyperparameter Tuning for Combining Multiple Distance Metrics
CRS-$v_{x}+v_{\mathrm{FID}}$ has four hyperparameters; $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$. We tuned these hyperparameters in two steps. In the first step, we fixed $\xi_{x}=\xi_{\mathrm{FID}}=1$ and searched for appropriate values of $w_{x}$ and $w_{\mathrm{FID}}$, subject to the constraint $w_{x}+w_{\mathrm{FID}}=1$. In the second step, we fixed $w_{x}$ and $w_{\mathrm{FID}}$, then adjusted $\xi_{x}$ and $\xi_{\mathrm{FID}}$.
The detailed results on LSUN Horse 256 $\times$ 256 are listed in Table 6. In the first step, adjusting $w_{x}$ and $w_{\mathrm{FID}}$ led to significant changes in FID at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. In the second step, the improvement from tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ was small at $\mathrm{NFE}=250$. In contrast, for $\mathrm{NFE}=50$, the tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ resulted in a substantial improvement in FID, indicating that the importance of hyperparameter tuning increases as NFE decreases.
Table 6: Detailed results of hyperparameter tuning on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model. DPM-Solver++(2M) was used as sampler. To conserve computational resources, FID score was calculated using 10K samples.
NFE Tuning step $w_{\mathrm{FID}}$ $x_{x}$ $\xi_{\mathrm{FID}}$ $\xi_{x}$ FID (10K)
250 1st 0.9 0.1 1.0 1.0 3.78
0.7 0.3 3.70
0.5 0.5 3.70
0.3 0.7 3.73
0.1 0.9 3.76
2nd 0.7 0.3 0.5 0.5 3.77
1.0 4.30
1.2 3.75
1.4 3.94
1.0 0.5 3.76
1.0 3.70
1.2 3.69
1.4 3.78
1.2 0.5 3.82
1.0 3.69
1.2 3.78
1.4 3.77
1.4 0.5 3.83
1.0 3.78
1.2 3.77
1.4 3.77
50 1st 0.9 0.1 1.0 1.0 6.13
0.7 0.3 4.42
0.5 0.5 4.38
0.3 0.7 4.78
0.1 0.9 7.03
2nd 0.5 0.5 0.5 0.5 7.42
1.0 7.44
1.2 7.48
1.4 7.79
1.0 0.5 4.85
1.0 4.38
1.2 4.32
1.4 4.27
1.2 0.5 4.57
1.0 4.09
1.2 3.96
1.4 3.93
1.4 0.5 4.51
1.0 3.95
1.2 3.85
1.4 3.82
The dependence of noise schedules on the hyperparameters is shown in Figs. 7, 7, and 7. Increasing either $w_{\mathrm{FID}}$ or $\xi_{\mathrm{FID}}$ amplifies the impact of $v_{\mathrm{FID}}(\alpha)$, resulting in more timesteps being allocated to the region where $\alpha\simeq 1$ (see Figs. 7 and 7). On the other hand, increasing either $w_{x}$ or $\xi_{x}$ strengthens the impact of $v_{x}(\alpha)$, causing more timesteps to be allocated to the region where $\alpha\simeq 0$ (see Figs. 7 and 7).
The hyperparameter settings for CRS-$v_{x}+v_{\mathrm{FID}}$ are listed in Table 7. Noise schedules $\alpha(t)$ obtained using CRS-$v_{x}+v_{\mathrm{FID}}$ are illustrated in Figs. 13, 13, 13, 13, 13, and 13.
Table 7: Hyperparameter settings of CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse 256 $\times$ 256, LSUN Bedroom 256 $\times$ 256, FFHQ 64 $\times$ 64, and CIFAR10 32 $\times$ 32.
Dataset Sampler NFE $w_{\mathrm{FID}}$ $w_{x}$ $\xi_{\mathrm{FID}}$ $\xi_{x}$
LSUN Horse 256 $\times$ 256 SDE-DPM-Solver++(2M) 250 0.9 0.1 1.4 1.0
50 0.7 0.3 1.4 1.2
PNDM 250 0.9 0.1 1.0 0.5
50 0.5 0.5 1.2 0.5
DPM-Solver++(2M) 250 0.7 0.3 1.0 1.2
50 0.5 0.5 1.4 1.4
LSUN Bedroom 256 $\times$ 256 DPM-Solver++(2M) 10 0.5 0.5 1.0 1.0
5 0.5 0.5 1.0 1.2
UniPC 10 0.5 0.5 1.0 1.2
5 0.3 0.7 1.0 1.0
FFHQ 64 $\times$ 64 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.4
20 0.5 0.5 1.6 1.4
10 0.5 0.5 1.6 1.5
iPNDM 50 0.5 0.5 1.2 1.5
20 0.5 0.5 1.4 1.4
10 0.5 0.5 1.5 1.4
CIFAR10 32 $\times$ 32 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.0
20 0.5 0.5 1.8 2.4
10 0.5 0.5 1.6 3.0
iPNDM 50 0.7 0.3 1.6 1.2
20 0.5 0.5 1.6 1.6
10 0.5 0.5 1.6 3.2
Figure 8: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with SDE-DPM-Solver++(2M).
Figure 9: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with PNDM.
Figure 10: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with DPM-Solver++(2M).
Figure 11: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with DPM-Solver++(2M).
Figure 12: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with UniPC.
Figure 13: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating FFHQ 64 $\times$ 64 with DPM-Solver++(2M).
Appendix E Implementation Details
E.1 Unet Model
The hyperparameters of the U-Net model are listed in Table 8. For latent-space diffusion models, we used the same values as the ImageNet 64 $\times$ 64 in ADM [52]. For pixel-space diffusion models, we used the same values as LSUN 256 $\times$ 256 in ADM [52].
Table 8: Hyperparameters of UNet model.
Latent-space diffusion models Pixel-space diffusion models
Resolution 64 $\times$ 64 256 $\times$ 256
Number of parameters 296M 552M
Channels 192 256
Depth 3 2
Channels multiple 1,2,3,4 1,1,2,2,4,4
Heads channels 64 64
Attention resolution 32,16,8 32,16,8
BigGAN up/downsample True True
Dropout 0.1 0.1
E.2 Hyperparameters for Training
The hyperparameters for training the diffusion models are summarized in Table 9. We saved ten checkpoints for each model. The FID score with 10K samples was evaluated for all checkpoints, and the results are reported with the checkpoint with the best FID score.
Table 9: Hyperparameters for training diffusion models. Same configuration was used for training latent-space and pixel-space diffusion models.
Item Value
EMA decay 0.9999
Optimizer Adam
$\beta_{1}$ 0.9
$\beta_{2}$ 0.999
Learning rate 1e-4
Batch size 32 $\times$ 8
E.3 Noise Schedule for Training and Sampling
The conventional noise schedules used in this study are summarized in Table 10.
When using a CRS-based training schedule, a common hyperparameter setting of $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.0)$ is used (details are described in Appendix C.2). When using a CRS-based sampling schedule, the default setting is $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.01)$. An exception was made when using CRS-$v_{x}+v_{\mathrm{FID}}$, where the hyperparameters are tuned according to the procedure described in Appendix D. For the experiments using pretrained EDMs, we used $(\alpha_{\max},\alpha_{\min})=(0.999998,0.0125)$, which is equivalent to the boundary condition of EDM: $(\sigma_{\min},\sigma_{\max})=(0.002,80)$. The boundary condition of EDM is transformed to $\alpha_{\max}$ and $\alpha_{\min}$ using Eq. (16).
Table 10: Conventional noise schedules evaluated in this study. $T_{s}$ denotes number of timesteps for sampling.
Name Phase Noise schedule
Linear Training
$\alpha_{\mathrm{linear}}(t)=\mathrm{LinearInterp1d}\left(\left\{\frac{i}{T}\right\}_{i=0,...,T},\left\{\alpha_{i}\right\}_{i=0,...,T}\right),$
$\alpha_{0}=1,$
$\alpha_{1\leq i\leq T}=\sqrt{1-\tilde{\beta}_{i}}\alpha_{i-1},$
$\tilde{\beta}_{i}=\tilde{\beta}_{\min}+\frac{\tilde{\beta}_{\max}-\tilde{\beta}_{\min}}{T-1}(i-1),$
$T=1000,\tilde{\beta}_{\min}=10^{-4},\tilde{\beta}_{\max}=0.02.$
Sampling $\left\{\alpha_{\mathrm{linear}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\}$
Shifted cosine Training
$\alpha_{\mathrm{shifted}}(t)=\sqrt{\mathrm{sigmoid}(\lambda(t))},$
$\lambda(t)=-2\log\tan\left(\frac{\pi t}{2}\right)+2\log\left(\frac{64}{d}\right),$
where $d$ is resolution in diffusion process.
Sampling
$\left\{\tilde{\alpha}_{\mathrm{shifted}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\},$
$\tilde{\alpha}_{\mathrm{shifted}}(t)=\alpha_{\min}+(\alpha_{\max}-\alpha_{\min})\alpha_{\mathrm{shifted}}(t),$
$\alpha_{\min}=0.01,\alpha_{\max}=1.0.$
VDM++ Training
$p(\lambda)\propto\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\omega(\lambda)\left\|\bm{\varepsilon}_{\theta}(\bm{x},\tilde{\alpha}(\lambda))-\bm{\varepsilon}\right\|_{2}^{2}\right],$
[…]
$\tilde{\omega}(\lambda)=\mathcal{N}(\lambda;\lambda_{\mu},\lambda_{\sigma}^{2})(e^{-\lambda}+c^{2}),$
$\mu_{\lambda}=2.4,\sigma_{\lambda}=2.4,c=0.5.$
Sampling Only for training.
EDM Training Not evaluated in this study.
Sampling
$\alpha_{0}=1,\alpha_{t\geq 1}=\alpha_{\mathrm{EDM}}\left(\frac{t}{T_{s}}\right),$
$\alpha_{\mathrm{EDM}}(t)=\sqrt{\mathrm{sigmoid}(\lambda(t))},$
$\lambda(t)=-2\log\sigma(t),$
$\sigma(t)=\left\{\sigma_{\max}^{\frac{1}{\rho}}+\left(\sigma_{\min}^{\frac{1}{\rho}}-\sigma_{\max}^{\frac{1}{\rho}}\right)(1-t)\right\}^{\rho},$
$\sigma_{\max}=80,\sigma_{\min}=0.002,\rho=7.$
Appendix F Detailed Results
We present the detailed experimental results and confirm the superiority of CRS for optimizing both training and sampling schedules.
F.1 Results on Pixel-Space Diffusion Models
LSUN Horse 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=250$ are listed in Table 11. To verify the generality with respect to samplers, we used three different samplers: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M). As shown in the rows shaded in gray, multiple evaluation metrics, including FID score, are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule, regardless of the sampler.
Table 12 presents the dependence on the sampling schedules at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. When using SDE-DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ improved upon all four evaluation metrics at $\mathrm{NFE}=250$, achieving a new state-of-the-art FID score. When using PNDM and DPM-Solver++(2M), there was little difference between sampling schedules at $\mathrm{NFE}=250$. However, at $\mathrm{NFE}=50$, the performance became more sensitive to the sampling schedule, and CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the best FID score.
LSUN Bedroom 256 $\times$ 256: Since only the FID score is reported in the main part of this study, we include sFID, precision, and recall in Table 13 for reference.
F.2 Results on Latent-Space Diffusion Models
LSUN Church 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 14. To verify the generality with respect to samplers, we used five different samplers: Stochastic DDIM, SDE-DPM-Solver++(2M), DDIM, PNDM, and DPM-Solver++(2M). As shown in the rows shaded in gray, FID scores are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule across all samplers except for SDE-DPM-Solver++(2M).
The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 15 and 16, respectively. CRS-$v_{x}$ or CRS-$v_{\varepsilon}$ achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
LSUN Bedroom 256 $\times$ 256: The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 17 and 18, respectively. CRS-$v_{x}$ is used for optimizing the training schedule. CRS achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=30$ and $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
FID scores evaluated at small NFEs are listed in Table 19. Compared with the results reported in DPM-Solver-v3 [26], the better FID scores are achieved by using CRS for optimizing both the training and sampling schedules at $\mathrm{NFE}=10$ and $\mathrm{NFE}=20$. FID, sFID, precision, and recall are listed in Table 20. CRS based noise schedules consistently outperform the linear schedule.
ImageNet 256 $\times$ 256: We demonstrated the superiority of CRS in class-conditional image generation. The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 21. As shown in the rows shaded in gray, CRS-$v_{\mathrm{FID}}$ was the best for optimizing training schedule for ImageNet, while CRS-$v_{x}$, which consistently showed strong performance on other datasets, was the second-best. The reason CRS-$v_{\mathrm{FID}}$ performed better on ImageNet may be attributed to the fact that the feature model used to compute FID is trained on ImageNet classification.
The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Tables 22 and 23, respectively. We used CRS-$v_{x}$ for optimizing training schedule, which outperformed conventional noise schedules with significant margin on ImageNet and exhibited the best performance across a wide range of datasets. CRS achieved the best FID score in all cases except when using PNDM at $\mathrm{NFE}=20$ and DPM-Solver++(2M) at $\mathrm{NFE}=30$, and multiple evaluation metrics are simultaneously improved in most cases.
F.3 Results on Pretrained EDM
FFHQ 64 $\times$ 64: The evaluation results are listed in Table 24. When using DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ outperforms both EDM and AYS. Using iPNDM further improves FID scores, and CRS-$v_{x}+v_{\mathrm{FID}}$ consistently performs better than EDM.
CIFAR10 32 $\times$ 32: The evaluation results on class-conditional image generation are listed in Table 25. CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperforms EDM, and significant performance gains are observed at small NFEs.
Appendix G Generated Samples
Generated images are shown in the following figures:
  • LSUN Horse 256 $\times$ 256 in the pixel-space diffusion model: Figs. 14 and 15.
  • LSUN Bedroom 256 $\times$ 256 in the pixel-space diffusion model: Figs. 16 and 17.
  • LSUN Church 256 $\times$ 256 in the latent-space diffusion model: Fig. 18.
  • LSUN Bedroom 256 $\times$ 256 in the latent-space diffusion model: Fig. 19.
  • ImageNet 256 $\times$ 256 in the latent-space diffusion model: Fig. 20.
All models are trained using CRS-$v_{x}$ for optimizing training schedule.
Table 11: Dependence on training and sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models at $\mathrm{NFE}=250$. To verify generality with respect to samplers, we used three different samplers. Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric. Note that "cosine" refers to shifted cosine in this table.
Training Sampling Metrics
Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
SDE-DPM- Linear Linear 2.90 6.82 0.66 0.56
Solver++(2M) Cosine 2.95 6.43 0.67 0.56
VDM++ 3.82 7.40 0.66 0.54
CRS-$v_{\varepsilon}$ 2.91 6.56 0.66 0.56
CRS-$v_{x}$ 2.73 6.40 0.67 0.56
CRS-$v_{x}$ Linear 2.73 6.40 0.67 0.56
Cosine 3.13 7.09 0.67 0.54
EDM 2.09 6.16 0.69 0.56
CRS-$v_{\varepsilon}$ 2.87 6.71 0.66 0.56
CRS-$v_{x}$ 5.46 8.34 0.62 0.52
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.03 6.06 0.69 0.56
PNDM Linear Linear 4.06 6.23 0.57 0.61
Cosine 2.36 5.59 0.63 0.60
VDM++ 70.18 36.78 0.20 0.53
CRS-$v_{\mathrm{FID}}$ 8.20 7.92 0.51 0.60
CRS-$v_{x}$ 2.09 5.43 0.65 0.58
CRS-$v_{x}$ Linear 2.09 5.43 0.65 0.58
Cosine 2.19 5.59 0.67 0.57
EDM 2.35 5.85 0.67 0.56
CRS-$v_{\mathrm{\varepsilon}}$ 2.11 5.58 0.66 0.58
CRS-$v_{x}$ 2.50 5.94 0.64 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.12 5.50 0.67 0.57
DPM- Linear Linear 3.17 5.98 0.60 0.61
Solver++(2M) Cosine 2.31 5.63 0.63 0.60
VDM++ 47.63 29.77 0.26 0.56
CRS-$v_{\varepsilon}$ 7.12 7.45 0.54 0.60
CRS-$v_{x}$ 2.08 5.47 0.65 0.59
CRS-$v_{x}$ Linear 2.08 5.47 0.65 0.59
Cosine 2.18 5.72 0.66 0.58
EDM 2.35 5.85 0.67 0.56
CRS-$v_{x}$ 2.12 5.65 0.66 0.58
CRS-$v_{\varepsilon}$ 2.87 6.15 0.63 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.19 5.56 0.67 0.57
Table 12: Dependence on sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
SDE-DPM- 250 Linear 2.73 6.40 0.67 0.56
Solver++(2M) Shifted cosine 3.13 7.09 0.67 0.54
EDM 2.09 6.16 0.69 0.56
CRS-$v_{\varepsilon}$ 2.87 6.71 0.66 0.56
CRS-$v_{x}$ 5.46 8.34 0.62 0.52
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.03 6.06 0.69 0.56
50 Linear 7.54 9.73 0.59 0.48
Shifted cosine 11.37 12.46 0.53 0.43
EDM 2.75 6.91 0.68 0.55
CRS-$v_{\varepsilon}$ 6.08 8.75 0.61 0.50
CRS-$v_{x}$ 12.99 12.59 0.50 0.41
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.84 8.16 0.67 0.52
PNDM 250 Linear 2.09 5.43 0.65 0.58
Shifted cosine 2.19 5.59 0.67 0.57
EDM 2.35 5.85 0.67 0.56
CRS-$v_{\varepsilon}$ 2.11 5.58 0.66 0.58
CRS-$v_{x}$ 2.50 5.94 0.64 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.12 5.50 0.67 0.57
50 Linear 2.50 5.76 0.62 0.59
Shifted cosine 3.92 6.77 0.60 0.57
EDM 2.36 5.87 0.67 0.56
CRS-$v_{\varepsilon}$ 2.46 5.97 0.64 0.58
CRS-$v_{x}$ 6.03 7.32 0.57 0.56
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.04 5.60 0.66 0.58
DPM- 250 Linear 2.08 5.47 0.65 0.59
Solver++(2M) Shifted Cosine 2.18 5.72 0.66 0.58
EDM 2.35 5.85 0.67 0.56
CRS-$v_{\varepsilon}$ 2.12 5.65 0.66 0.58
CRS-$v_{x}$ 2.87 6.15 0.63 0.58
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.19 5.56 0.67 0.57
50 Linear 3.17 5.98 0.60 0.59
Shifted cosine 5.27 7.12 0.57 0.56
EDM 2.43 5.87 0.67 0.56
CRS-$v_{\varepsilon}$ 3.10 6.18 0.62 0.58
CRS-$v_{x}$ 7.44 7.83 0.55 0.55
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.26 5.90 0.66 0.57
Table 13: Evaluation results on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion model. CRS-$v_{x}$ was used for optimizing training schedule.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DPM-Solver++(2M) 10 EDM 4.73 7.75 0.53 0.51
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.88 8.67 0.49 0.52
5 EDM 23.72 19.24 0.20 0.28
CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 17.49 0.28 0.41
UniPC 10 EDM 4.94 8.20 0.53 0.40
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.30 6.60 0.55 0.53
5 EDM 82.52 33.35 0.06 0.08
CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 11.26 0.21 0.44
Table 14: Dependence on training and sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$. To verify generality with respect to samplers, we used five different samplers. Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric.
Training Sampling Metrics
Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM Linear Linear 10.73 18.31 0.57 0.33
VDM++ 11.05 17.33 0.56 0.35
CRS-$v_{\mathrm{FID}}$ 10.85 18.77 0.57 0.34
CRS-$v_{\varepsilon}$ 10.43 17.81 0.57 0.34
CRS-$v_{x}$ 9.82 16.39 0.58 0.34
CRS-$v_{x}$ Linear 9.82 16.39 0.58 0.34
EDM 21.73 29.33 0.43 0.22
CRS-$v_{\mathrm{FID}}$ 7.36 13.04 0.60 0.39
CRS-$v_{\varepsilon}$ 7.53 13.73 0.61 0.38
CRS-$v_{x}$ 6.77 12.57 0.63 0.40
SDE-DPM- Linear Linear 3.49 10.19 0.63 0.57
Solver++(2M) VDM++ 3.33 9.79 0.63 0.57
CRS-$v_{\mathrm{FID}}$ 3.50 10.69 0.63 0.57
CRS-$v_{\varepsilon}$ 3.57 10.26 0.64 0.56
CRS-$v_{x}$ 3.38 10.09 0.64 0.55
CRS-$v_{x}$ Linear 3.38 10.09 0.64 0.55
EDM 4.20 11.69 0.63 0.54
CRS-$v_{\mathrm{FID}}$ 3.61 9.51 0.61 0.57
CRS-$v_{\varepsilon}$ 3.19 9.56 0.62 0.59
CRS-$v_{x}$ 3.20 9.14 0.62 0.58
DDIM Linear Linear 6.18 13.16 0.59 0.51
VDM++ 6.50 12.67 0.56 0.54
CRS-$v_{\mathrm{FID}}$ 6.06 13.69 0.59 0.51
CRS-$v_{\varepsilon}$ 6.05 12.49 0.58 0.50
CRS-$v_{x}$ 5.67 12.33 0.59 0.51
CRS-$v_{x}$ Linear 5.67 12.33 0.59 0.51
EDM 9.15 16.86 0.56 0.44
CRS-$v_{\mathrm{FID}}$ 5.35 10.70 0.57 0.53
CRS-$v_{\varepsilon}$ 4.95 11.01 0.59 0.53
CRS-$v_{x}$ 4.81 10.79 0.59 0.55
PNDM Linear Linear 3.74 10.48 0.60 0.58
VDM++ 3.88 10.20 0.59 0.59
CRS-$v_{\mathrm{FID}}$ 3.59 10.61 0.61 0.59
CRS-$v_{\varepsilon}$ 3.71 10.36 0.60 0.59
CRS-$v_{x}$ 3.56 10.37 0.61 0.58
CRS-$v_{x}$ Linear 3.56 10.37 0.61 0.58
EDM 3.88 10.76 0.60 0.57
CRS-$v_{\mathrm{FID}}$ 3.59 9.69 0.59 0.61
CRS-$v_{\varepsilon}$ 3.48 9.79 0.59 0.59
CRS-$v_{x}$ 3.62 10.15 0.59 0.61
DPM- Linear Linear 3.82 10.28 0.60 0.58
Solver++(2M) VDM++ 4.00 10.07 0.58 0.59
CRS-$v_{\mathrm{FID}}$ 3.78 10.51 0.60 0.58
CRS-$v_{\varepsilon}$ 3.87 10.30 0.60 0.59
CRS-$v_{x}$ 3.69 10.23 0.60 0.59
CRS-$v_{x}$ Linear 3.69 10.23 0.60 0.59
EDM 4.31 11.31 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.98 9.61 0.58 0.60
CRS-$v_{\varepsilon}$ 3.63 9.91 0.59 0.60
CRS-$v_{x}$ 3.59 9.51 0.59 0.60
Table 15: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM 50 Linear 6.23 11.94 0.65 0.43
EDM 10.05 17.14 0.60 0.34
CRS-$v_{\mathrm{FID}}$ 5.17 10.49 0.65 0.44
CRS-$v_{\varepsilon}$ 5.26 10.94 0.65 0.44
CRS-$v_{x}$ 4.82 11.91 0.62 0.51
30 Linear 9.82 16.39 0.58 0.34
EDM 21.73 29.33 0.43 0.22
CRS-$v_{\mathrm{FID}}$ 7.36 13.04 0.60 0.39
CRS-$v_{\varepsilon}$ 7.53 13.73 0.61 0.38
CRS-$v_{x}$ 6.77 12.57 0.63 0.40
20 Linear 18.24 25.36 0.45 0.24
EDM 48.46 52.53 0.25 0.11
CRS-$v_{\mathrm{FID}}$ 12.42 18.38 0.51 0.29
CRS-$v_{\varepsilon}$ 12.53 19.27 0.52 0.29
CRS-$v_{x}$ 10.66 16.54 0.55 0.32
SDE-DPM- 50 Linear 3.42 10.58 0.63 0.57
Solver++(2M) EDM 3.57 10.93 0.64 0.57
CRS-$v_{\mathrm{FID}}$ 3.32 9.19 0.62 0.58
CRS-$v_{\varepsilon}$ 3.14 9.65 0.63 0.58
CRS-$v_{x}$ 3.09 9.34 0.63 0.59
30 Linear 3.38 10.09 0.64 0.55
EDM 4.20 11.69 0.63 0.54
CRS-$v_{\mathrm{FID}}$ 3.61 9.51 0.61 0.57
CRS-$v_{\varepsilon}$ 3.19 9.56 0.62 0.59
CRS-$v_{x}$ 3.20 9.14 0.62 0.58
20 Linear 3.55 10.37 0.63 0.56
EDM 5.92 13.79 0.62 0.52
CRS-$v_{\mathrm{FID}}$ 4.39 9.33 0.63 0.48
CRS-$v_{\varepsilon}$ 3.58 9.93 0.61 0.57
CRS-$v_{x}$ 3.57 9.55 0.62 0.56
Table 16: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DDIM 50 Linear 4.52 11.03 0.60 0.54
EDM 5.95 13.37 0.59 0.50
CRS-$v_{\mathrm{FID}}$ 4.28 10.07 0.59 0.56
CRS-$v_{\varepsilon}$ 4.17 10.66 0.60 0.56
CRS-$v_{x}$ 4.01 10.10 0.60 0.57
30 Linear 5.67 12.33 0.59 0.51
EDM 9.15 16.86 0.56 0.44
CRS-$v_{\mathrm{FID}}$ 5.35 10.70 0.57 0.53
CRS-$v_{\varepsilon}$ 4.95 11.01 0.59 0.53
CRS-$v_{x}$ 4.81 10.79 0.59 0.55
20 Linear 7.85 14.66 0.56 0.47
EDM 15.74 22.86 0.48 0.34
CRS-$v_{\mathrm{FID}}$ 7.47 12.74 0.54 0.48
CRS-$v_{\varepsilon}$ 6.49 12.55 0.56 0.49
CRS-$v_{x}$ 6.09 11.72 0.57 0.50
PNDM 50 Linear 3.52 10.42 0.61 0.58
EDM 3.71 10.59 0.60 0.57
CRS-$v_{\mathrm{FID}}$ 3.49 9.95 0.60 0.61
CRS-$v_{\varepsilon}$ 3.48 10.18 0.60 0.59
CRS-$v_{x}$ 3.48 9.92 0.60 0.60
30 Linear 3.56 10.37 0.61 0.58
EDM 3.88 10.76 0.60 0.57
CRS-$v_{\mathrm{FID}}$ 3.59 9.69 0.59 0.61
CRS-$v_{\varepsilon}$ 3.48 9.79 0.59 0.59
CRS-$v_{x}$ 3.62 10.15 0.59 0.61
20 Linear 3.74 10.24 0.60 0.59
EDM 4.57 11.49 0.59 0.56
CRS-$v_{\mathrm{FID}}$ 4.06 9.54 0.57 0.61
CRS-$v_{\varepsilon}$ 3.64 9.98 0.59 0.61
CRS-$v_{x}$ 3.86 9.54 0.62 0.52
DPM- 50 Linear 3.62 10.44 0.61 0.58
Solver++(2M) EDM 3.90 10.85 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.59 9.68 0.59 0.60
CRS-$v_{\varepsilon}$ 3.46 9.82 0.60 0.59
CRS-$v_{x}$ 3.43 9.45 0.60 0.59
30 Linear 3.69 10.23 0.60 0.59
EDM 4.31 11.31 0.60 0.56
CRS-$v_{\mathrm{FID}}$ 3.98 9.61 0.58 0.60
CRS-$v_{\varepsilon}$ 3.63 9.91 0.59 0.60
CRS-$v_{x}$ 3.59 9.51 0.59 0.60
20 Linear 3.92 10.03 0.59 0.59
EDM 5.25 12.19 0.59 0.55
CRS-$v_{\mathrm{FID}}$ 4.95 9.92 0.56 0.58
CRS-$v_{\varepsilon}$ 3.89 9.73 0.58 0.60
CRS-$v_{x}$ 4.00 9.44 0.58 0.59
Table 17: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM 50 Linear 3.95 8.98 0.56 0.48
EDM 7.43 16.05 0.54 0.38
CRS-$v_{\mathrm{FID}}$ 3.09 7.17 0.57 0.50
CRS-$v_{\varepsilon}$ 3.25 7.81 0.57 0.49
CRS-$v_{x}$ 3.08 7.20 0.58 0.50
30 Linear 7.28 15.70 0.53 0.38
EDM 17.03 33.60 0.40 0.26
CRS-$v_{\mathrm{FID}}$ 5.17 10.62 0.57 0.43
CRS-$v_{\varepsilon}$ 5.58 12.33 0.56 0.42
CRS-$v_{x}$ 4.78 9.85 0.55 0.45
20 Linear 14.53 29.25 0.42 0.28
EDM 38.01 68.62 0.21 0.13
CRS-$v_{\mathrm{FID}}$ 9.18 16.67 0.50 0.35
CRS-$v_{\varepsilon}$ 9.70 19.40 0.49 0.34
CRS-$v_{x}$ 8.11 15.96 0.51 0.36
SDE-DPM- 50 Linear 2.47 6.44 0.57 0.54
Solver++(2M) EDM 2.48 6.32 0.57 0.54
CRS-$v_{\mathrm{FID}}$ 2.21 5.87 0.57 0.54
CRS-$v_{\varepsilon}$ 2.22 5.76 0.57 0.55
CRS-$v_{x}$ 2.23 5.67 0.57 0.54
30 Linear 2.45 6.70 0.56 0.55
EDM 2.90 7.41 0.57 0.52
CRS-$v_{\mathrm{FID}}$ 2.46 6.19 0.57 0.54
CRS-$v_{\varepsilon}$ 2.50 5.93 0.57 0.53
CRS-$v_{x}$ 2.48 5.73 0.56 0.54
20 Linear 2.78 6.79 0.56 0.53
EDM 4.00 9.57 0.55 0.50
CRS-$v_{\mathrm{FID}}$ 3.19 6.70 0.55 0.53
CRS-$v_{\varepsilon}$ 2.87 6.22 0.57 0.53
CRS-$v_{x}$ 3.07 6.34 0.56 0.52
Table 18: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DDIM 50 Linear 3.02 6.59 0.53 0.54
EDM 4.17 7.64 0.52 0.51
CRS-$v_{\mathrm{FID}}$ 2.66 6.03 0.54 0.55
CRS-$v_{\varepsilon}$ 2.71 6.03 0.54 0.55
CRS-$v_{x}$ 2.62 5.86 0.54 0.55
30 Linear 3.76 7.50 0.53 0.51
EDM 6.81 9.98 0.47 0.46
CRS-$v_{\mathrm{FID}}$ 3.23 6.61 0.53 0.53
CRS-$v_{\varepsilon}$ 3.22 6.65 0.53 0.53
CRS-$v_{x}$ 3.10 6.43 0.54 0.53
20 Linear 5.38 9.27 0.49 0.48
EDM 12.57 14.55 0.38 0.40
CRS-$v_{\mathrm{FID}}$ 4.39 7.86 0.51 0.49
CRS-$v_{\varepsilon}$ 4.28 7.97 0.51 0.49
CRS-$v_{x}$ 4.10 7.54 0.51 0.51
PNDM 50 Linear 2.51 6.02 0.54 0.56
EDM 2.64 6.16 0.53 0.57
CRS-$v_{\mathrm{FID}}$ 2.50 5.66 0.54 0.57
CRS-$v_{\varepsilon}$ 2.44 5.73 0.54 0.57
CRS-$v_{x}$ 2.45 5.61 0.54 0.56
30 Linear 2.57 6.03 0.54 0.56
EDM 2.93 6.45 0.52 0.57
CRS-$v_{\mathrm{FID}}$ 2.64 5.59 0.53 0.57
CRS-$v_{\varepsilon}$ 2.48 5.58 0.54 0.56
CRS-$v_{x}$ 2.65 5.53 0.54 0.57
20 Linear 2.77 6.13 0.53 0.57
EDM 3.76 7.25 0.52 0.54
CRS-$v_{\mathrm{FID}}$ 3.02 5.66 0.52 0.57
CRS-$v_{\varepsilon}$ 2.71 5.53 0.54 0.56
CRS-$v_{x}$ 3.63 6.23 0.61 0.57
DPM- 50 Linear 2.59 6.15 0.54 0.56
Solver++(2M) EDM 2.86 6.44 0.52 0.57
CRS-$v_{\mathrm{FID}}$ 2.46 5.63 0.54 0.56
CRS-$v_{\varepsilon}$ 2.46 5.63 0.54 0.56
CRS-$v_{x}$ 2.37 5.42 0.54 0.56
30 Linear 2.68 6.18 0.54 0.56
EDM 3.40 7.09 0.51 0.56
CRS-$v_{\mathrm{FID}}$ 2.60 5.63 0.54 0.56
CRS-$v_{\varepsilon}$ 2.60 5.53 0.54 0.56
CRS-$v_{x}$ 2.63 5.45 0.54 0.55
20 Linear 2.89 6.24 0.53 0.56
EDM 4.14 8.07 0.52 0.53
CRS-$v_{\mathrm{FID}}$ 3.03 5.81 0.52 0.56
CRS-$v_{\varepsilon}$ 2.88 5.59 0.53 0.56
CRS-$v_{x}$ 3.00 5.63 0.54 0.55
Table 19: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs. Bold values indicate best FID scores at each NFE. Results for rows other than "Ours" are taken from previously published work [26].
Sampling FID $\downarrow$
Model Sampler schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$ $\mathrm{NFE}=20$
LDM [57] DPM-Solver++ Linear 18.59 3.63 3.16
UniPC 12.24 3.56 3.07
DPM-Solver-v3 [26] 7.54 3.16 3.05
Ours iPNDM Linear 10.81 3.71 2.64
CRS-$v_{\varepsilon}$ 7.88 3.12 2.48
CRS-$v_{x}$ 8.07 3.45 2.50
Table 20: Detailed results evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
iPNDM 20 Linear 2.64 6.15 0.53 0.57
CRS-$v_{\varepsilon}$ 2.48 5.48 0.54 0.56
CRS-$v_{x}$ 2.50 5.41 0.55 0.56
10 Linear 3.71 6.88 0.52 0.54
CRS-$v_{\varepsilon}$ 3.12 5.69 0.52 0.56
CRS-$v_{x}$ 3.45 6.15 0.54 0.53
5 Linear 10.81 13.82 0.42 0.49
CRS-$v_{\varepsilon}$ 7.88 9.84 0.45 0.50
CRS-$v_{x}$ 8.07 10.43 0.45 0.48
Table 21: Dependence on training and sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$. To verify generality with respect to samplers, we used five different samplers. Rows shaded in gray represent dependence on training schedule, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric.
Training Sampling Metrics
Sampler schedule schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM Linear Linear 27.31 21.39 0.46 0.55
VDM++ 29.00 17.67 0.46 0.57
CRS-$v_{\mathrm{FID}}$ 25.50 21.41 0.48 0.56
CRS-$v_{\varepsilon}$ 26.97 20.71 0.47 0.56
CRS-$v_{x}$ 25.83 19.92 0.47 0.57
CRS-$v_{x}$ Linear 25.83 19.92 0.47 0.57
EDM 49.72 50.08 0.31 0.51
CRS-$v_{\mathrm{FID}}$ 20.85 13.01 0.52 0.56
CRS-$v_{\varepsilon}$ 20.43 13.14 0.53 0.58
CRS-$v_{x}$ 17.88 10.14 0.55 0.59
SDE-DPM- Linear Linear 11.12 5.68 0.62 0.62
Solver++(2M) VDM++ 29.00 17.67 0.46 0.57
CRS-$v_{\mathrm{FID}}$ 10.22 5.81 0.62 0.64
CRS-$v_{\varepsilon}$ 11.18 5.84 0.62 0.63
CRS-$v_{x}$ 10.25 5.64 0.62 0.64
CRS-$v_{x}$ Linear 10.25 5.64 0.62 0.64
EDM 11.52 6.13 0.60 0.64
CRS-$v_{\mathrm{FID}}$ 10.49 5.39 0.63 0.63
CRS-$v_{\varepsilon}$ 10.24 5.22 0.63 0.63
CRS-$v_{x}$ 10.56 5.32 0.63 0.63
DDIM Linear Linear 17.33 6.79 0.55 0.62
VDM++ 19.53 5.97 0.54 0.64
CRS-$v_{\mathrm{FID}}$ 15.52 6.95 0.56 0.63
CRS-$v_{\varepsilon}$ 17.02 6.94 0.55 0.63
CRS-$v_{x}$ 16.28 6.27 0.56 0.64
CRS-$v_{x}$ Linear 16.28 6.27 0.56 0.64
EDM 22.17 9.22 0.49 0.62
CRS-$v_{\mathrm{FID}}$ 15.51 6.72 0.56 0.63
CRS-$v_{\varepsilon}$ 15.26 5.79 0.57 0.64
CRS-$v_{x}$ 15.28 5.67 0.57 0.64
PNDM Linear Linear 13.21 5.35 0.59 0.65
VDM++ 16.01 6.31 0.57 0.66
CRS-$v_{\mathrm{FID}}$ 11.42 5.31 0.59 0.66
CRS-$v_{\varepsilon}$ 12.69 5.45 0.59 0.65
CRS-$v_{x}$ 12.29 5.38 0.59 0.66
CRS-$v_{x}$ Linear 12.29 5.38 0.59 0.66
EDM 12.86 5.62 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 12.18 5.21 0.59 0.65
CRS-$v_{\varepsilon}$ 12.57 5.24 0.59 0.66
CRS-$v_{x}$ 12.65 5.38 0.59 0.66
DPM- Linear Linear 13.87 5.41 0.59 0.64
Solver++(2M) VDM++ 16.97 6.45 0.56 0.66
CRS-$v_{\mathrm{FID}}$ 12.09 5.40 0.59 0.65
CRS-$v_{\varepsilon}$ 13.41 5.54 0.58 0.65
CRS-$v_{x}$ 12.92 5.46 0.59 0.66
CRS-$v_{x}$ Linear 12.92 5.46 0.59 0.66
EDM 13.67 5.79 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 13.02 5.19 0.59 0.65
CRS-$v_{\varepsilon}$ 13.02 5.17 0.59 0.65
CRS-$v_{x}$ 13.33 5.13 0.59 0.65
Table 22: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Stochastic DDIM 50 Linear 15.61 8.97 0.57 0.60
EDM 25.66 19.17 0.46 0.57
CRS-$v_{\mathrm{FID}}$ 13.67 6.86 0.60 0.60
CRS-$v_{\varepsilon}$ 13.44 7.56 0.59 0.61
CRS-$v_{x}$ 13.05 6.45 0.60 0.61
30 Linear 25.83 19.92 0.47 0.57
EDM 49.72 50.08 0.31 0.51
CRS-$v_{\mathrm{FID}}$ 20.85 13.01 0.52 0.56
CRS-$v_{\varepsilon}$ 20.43 13.14 0.53 0.58
CRS-$v_{x}$ 17.88 10.14 0.55 0.59
20 Linear 44.15 42.96 0.34 0.53
EDM 78.05 99.25 0.20 0.44
CRS-$v_{\mathrm{FID}}$ 30.28 24.26 0.44 0.54
CRS-$v_{\varepsilon}$ 31.86 26.34 0.43 0.55
CRS-$v_{x}$ 27.20 20.09 0.47 0.56
SDE-DPM- 50 Linear 9.96 5.62 0.63 0.64
DPM-Solver++(2M) EDM 9.72 5.35 0.63 0.64
CRS-$v_{\mathrm{FID}}$ 10.26 5.43 0.63 0.63
CRS-$v_{\varepsilon}$ 9.68 5.33 0.63 0.64
CRS-$v_{x}$ 10.11 5.32 0.63 0.64
30 Linear 10.25 5.64 0.62 0.64
EDM 11.52 6.13 0.60 0.64
CRS-$v_{\mathrm{FID}}$ 10.49 5.39 0.63 0.63
CRS-$v_{\varepsilon}$ 10.24 5.22 0.63 0.63
CRS-$v_{x}$ 10.56 5.32 0.63 0.63
20 Linear 11.47 5.81 0.61 0.63
EDM 16.26 9.38 0.55 0.63
CRS-$v_{\mathrm{FID}}$ 11.10 5.50 0.63 0.62
CRS-$v_{\varepsilon}$ 10.83 5.36 0.62 0.63
CRS-$v_{x}$ 11.64 5.26 0.63 0.62
Table 23: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
Sampling Metrics
Sampler NFE schedule FID $\downarrow$ sFID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DDIM 50 Linear 13.91 5.37 0.58 0.65
EDM 16.25 6.24 0.55 0.65
CRS-$v_{\mathrm{FID}}$ 13.45 5.36 0.58 0.64
CRS-$v_{\varepsilon}$ 13.64 5.20 0.58 0.65
CRS-$v_{x}$ 13.75 5.12 0.59 0.65
30 Linear 16.28 6.27 0.56 0.64
EDM 22.17 9.22 0.49 0.62
CRS-$v_{\mathrm{FID}}$ 15.51 6.72 0.56 0.63
CRS-$v_{\varepsilon}$ 15.26 5.79 0.57 0.64
CRS-$v_{x}$ 15.28 5.67 0.57 0.64
20 Linear 20.49 8.82 0.51 0.62
EDM 33.08 15.78 0.40 0.60
CRS-$v_{\mathrm{FID}}$ 18.43 7.52 0.54 0.63
CRS-$v_{\varepsilon}$ 18.30 7.57 0.54 0.63
CRS-$v_{x}$ 18.16 7.05 0.55 0.62
PNDM 50 Linear 12.10 5.32 0.59 0.66
EDM 12.53 5.49 0.59 0.66
CRS-$v_{\mathrm{FID}}$ 11.86 5.20 0.59 0.65
CRS-$v_{\varepsilon}$ 12.35 5.28 0.59 0.66
CRS-$v_{x}$ 12.44 5.28 0.59 0.66
30 Linear 12.29 5.38 0.59 0.66
EDM 12.86 5.62 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 12.18 5.21 0.59 0.65
CRS-$v_{\varepsilon}$ 12.57 5.24 0.59 0.66
CRS-$v_{x}$ 12.65 5.38 0.59 0.66
20 Linear 12.83 5.52 0.59 0.66
EDM 13.99 5.99 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 12.97 5.38 0.58 0.66
CRS-$v_{\varepsilon}$ 12.92 5.30 0.59 0.65
CRS-$v_{x}$ 13.63 6.10 0.57 0.66
DPM- 50 Linear 12.42 5.41 0.59 0.66
Solver++(2M) EDM 12.90 5.57 0.59 0.66
CRS-$v_{\mathrm{FID}}$ 12.27 5.15 0.59 0.65
CRS-$v_{\varepsilon}$ 12.64 5.23 0.59 0.66
CRS-$v_{x}$ 12.75 5.21 0.59 0.65
30 Linear 12.92 5.46 0.59 0.66
EDM 13.67 5.79 0.58 0.66
CRS-$v_{\mathrm{FID}}$ 13.02 5.19 0.59 0.65
CRS-$v_{\varepsilon}$ 13.02 5.17 0.59 0.65
CRS-$v_{x}$ 13.33 5.13 0.59 0.65
20 Linear 13.71 5.45 0.58 0.65
EDM 15.31 6.23 0.57 0.66
CRS-$v_{\mathrm{FID}}$ 14.10 5.25 0.59 0.65
CRS-$v_{\varepsilon}$ 13.70 5.13 0.59 0.65
CRS-$v_{x}$ 14.16 5.09 0.59 0.64
Table 24: Evaluation results on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM. Bold values indicate best sampling schedules for each combination of sampler and NFE. FID scores on AYS is taken from their original paper [16].
Sampling FID $\downarrow$
Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) EDM 5.57 3.02 2.58
AYS 5.43 3.29 2.62
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
iPNDM EDM 4.60 2.62 2.49
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.87 2.53 2.44
Table 25: Evaluation results for CIFAR10 32 $\times$ 32 on class-conditional image generation using the pretrained model provided in EDM. Bold values indicate best sampling schedules for each combination of sampler and NFE.
Sampling FID $\downarrow$
Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) EDM 4.12 2.20 1.94
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.03 2.07 1.89
iPNDM EDM 2.51 1.93 1.89
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.33 1.92 1.89
Figure 14: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=250$ ($\mathrm{FID}=2.03$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. SDE-DPM-Solver++(2M) was used as sampler.
Figure 15: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=50$ ($\mathrm{FID}=2.04$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. PNDM was used as sampler.
Figure 16: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=10$ ($\mathrm{FID}=3.30$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. UniPC was used as sampler.
Figure 17: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=5$ ($\mathrm{FID}=14.02$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. DPM-Solver++(2M) was used as sampler.
Figure 18: Samples of LSUN Church 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=3.89$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. DPM-Solver++(2M) was used as sampler.
Figure 19: Samples of LSUN Bedroom 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=2.71$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. PNDM was used as sampler.
Figure 20: Samples of ImageNet 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ with class condition ($\mathrm{FID}=10.83$). Six classes were randomly selected, and each row corresponds to one class. We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. SDE-DPM-Solver++(2M) was used as sampler.